# SPECIAL RELATIVITY

### v9

@@ 13.8.1 Phenomenology and Relativistic Extensions
In order fully to understand the meaning of the embedding of the Newton’s dynamics law in the Minkowski space-time, one must keep in mind that, above, we characterized Newton’s equation in the Lorentz comoving frame as a phenomenological law. The microscopic interpretation of the inertial mass of a particle is not given. In other words, it is generally accepted that Newton’s second law is a phenomenological law and the rest mass is introduced in an ad hoc manner. The system of coordinates in which the equations of Newton’s mechanics are valid can be defined as Lorentz rest frame. The relativistic generalization of the Newton’s second law to any Lorentz frame permits us to make correct predictions.
We are in the position to formulate the following general statement: any phenomenological law, which is valid in the Lorentz rest frame, can be embedded in the four dimensional space-time only by using Lorentz coordinatization (i.e. Einstein synchronization convention). Suppose we do not know why a muon disintegrates, but we know the law of decay in the Lorentz rest frame. This law would then be a phenomenological law. The relativistic generalization of this law to any Lorentz frame allows us to make a prediction on the average distance traveled by the muon. In contrast, in the non covariant (3+1) space and time approach there is no time dilation effect, since for Galilean transformations the time scales do not change. Therefore, in the (3+1) non covariant approach, there is no kinematics correction factor $\gamma$ to the travel distance of relativistically moving muons. The two approaches give, in fact, a different result for the travel-distance, which must be, however, convention-invariant. This glaring conflict between results of covariant and non covariant approaches can be explained as follows: it is a dynamical line of arguments that explains this paradoxical situation with the relativistic $\gamma$ factor. In fact, there is a machinery behind the muon disintegration. Its origin is explained in the framework of the Lorentz-covariant quantum field theory. In the microscopic approach to muon disintegration, Einstein and absolute time synchronization conventions give the same result for such convention-invariant observables like the average travel distance, and it does not matter which transformation (Galilean or Lorentz) is used.
13.8.2 The Relativistic Mass
In the non covariant (3+1) space and time approach, there is no time dilation nor length contraction, because for Galilean transformations time and spatial coordinates scales do not change. Moreover, it can easily be verified that Newton’s second law keeps its form under Galilean transformations. Therefore, in the (3+1) non covariant approach, there is no kinematics correction factor $\gamma$ to the mass in Newton’s second law. However, in contrast to kinematics effects like time dilation and length contraction, the correction factor $\gamma$ to the mass in the Newton’s second law has direct objective meaning. In fact, if we assign space-time coordinates to the lab frame using the absolute time convention, the equation of motion is still given by Newton’s second law corrected for the relativistic dependence of momentum on velocity even though, as just stated, it has no kinematical origin. Understanding this result of the theory of relativity is similar to understanding previously discussed results: at first we use Lorentz coordinates and later the (3+1) non covariant approach in terms of a microscopic interpretation that must be consistent with the principle of relativity.
It is well-known from classical electrodynamics that the electromagnetic field of an electron carries a momentum proportional to its velocity for $v\ll c$, while for an arbitrary velocity $v$, the momentum is altered by the relativistic $\gamma$ factor in the case when the absolute time convention is used. Many attempts have been made to explain the electron mass as fully originating from electromagnetic fields. However, these attempts have failed. In fact, it is impossible to have a stationary non-neutral charge distribution held together by purely electromagnetic forces. In other words, mass and momentum of an electron cannot be completely electromagnetic in origin and in order to grant stability there is a necessity for compensating electromagnetic forces with non electromagnetic fields. From this viewpoint, Newton’s second law is an empirical phenomenological law where the relativistic correction factor $\gamma$ to the mass is introduced in an ad hoc manner.
From a microscopic viewpoint, today accepted explanation of how structureless particles like leptons and quarks acquire mass is based on the coupling to the Higgs field, the Higgs boson having been recently experimentally observed at the LHC. This mechanism can be invoked to explain Newton’s second law from a microscopic viewpoint even for structureless particles like electrons. However, at larger scales, an interesting and intuitive concept of the origin of physical inertia is illustrated, without recurring to the Higgs field, by results of Quantum Chromodynamics (QCD) for protons and neutrons, which are not elementary and are composed of quarks and gluon fields. If an initial, unperturbed nuclear configuration is disturbed, the gluon field generates forces that tend to restore this unperturbed configuration. It is the distortion of the nuclear field that gives rise to the force in opposition to the one producing it, in analogy to the electromagnetic case. But in contrast to the electromagnetic model of an electron, the QCD model of a nucleon is stable, and other compensation fields are not needed. Now, the gluon field mass can be computed from the total energy (or momentum) stored in the field, and it turns out that the QCD version in which quark masses are taken as zero provides a remarkably good approximation to reality. Since this version of QCD is a theory whose basic building blocks have zero mass, the most of the mass of ordinary matter (more than 90 percent) arises from pure field energy. In other words, the mass of a nucleon can be explained almost entirely from a microscopic viewpoint, which automatically provides a microscopic explanation of Newton’s second law of motion. In order to predict, on dynamical grounds, the inertial mass of a relativistically moving nucleon one does not need to have access to the detailed dynamics of strong interactions. It is enough to assume Lorentz covariance (i.e. Lorentz form-invariance of field equations) of the complete QCD dynamics involved in nucleon mass calculations.
The previous discussion, results in a most general statement: it is enough to assume Lorentz covariance of the quantum field theory involved in micro-particle (elementary or not elementary) mass calculations in order to obtain the same result for the relativistic mass correction from the two synchronization conventions discussed here, and it does not matter which transformation (Galilean or Lorentz) is used.
13.8.3 What does Space-Time Geometry Explain?
It is important to stress at this point that the dynamical line of argument discussed here explains what the Minkowski geometry physically means. The pseudo-Euclidean geometric structure of space-time is only an interpretation of the behavior of the dynamical matter fields in the view of different observers, which is an observable, empirical fact. It should be clear that the relativistic properties of the dynamical matter fields are fundamental, while the geometric structure is not. Dynamics, based on the field equations, is actually hidden in the language of kinematics. The Lorentz covariance of the equations that govern the fundamental interactions of nature is an empirical fact, while the postulation of the pseudo-Euclidean geometry of space-time is a mathematical interpretation of it that yields the laws of relativistic kinematics: at a fundamental level this postulate is, however, based on the way fields behave dynamically.
13.9 Bibliography and Notes
1. Let us try to get a better understanding of the geometric restatement of Newton’s second law. To derive the covariant form of relativistic dynamics, we should embed the three-dimensional vector relation $md\vec{v}/dt=\vec{f}$ into the four-dimensional geometry of Minkowski space [52]. The idea of embedding is based on the principle of relativity i.e. on the fact that the usual Newton’s second law can always be used in any Lorentz frame where the particle, whose motion we want to describe, is at rest. In other words, if an instantaneously comoving Lorentz frame is given at some instant, one can precisely predict the evolution of the particle in this frame during an infinitesimal time interval. In geometric language, the Newton law is strict on a hyperplane perpendicular to the world line. However, the hyperplane tilts together with its normal $u_{\mu}$ as one moves along the world line. For the embedding we need an operator $\hat{P}_{\perp}$ that continually projects vectors of Minkowski space on hyperplanes perpendicular to world line. The desired operator is $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [52]. In the instantaneously comoving frame one can unambiguously construct a four-force $K_{\mu}=[0,\vec{f}]$. Then, in an arbitrary Lorentz frame, the components $K_{\mu}$ can be found through the appropriate Lorentz transformation. In the rest frame obviously $u_{\mu}K^{\mu}=0$. It follows that, since $u_{\mu}K^{\mu}$ is an invariant, the four-force $K_{\mu}$ is perpendicular to the four-velocity $u_{\mu}$ in any Lorentz frame. The desired embedding of Newton’s second law in hyperplanes perpendicular to the world line is found by imposing $(\hat{P}_{\perp})_{\mu\nu}(mdu_{\nu}/d\tau-K_{\nu})=0$. This is a tensor equation in Minkowski space-time that relates geometric objects and does not need coordinates to be expressed. The evolution of a particle can be described in terms of world line $\sigma(\tau)$, and the 4-velocity by $u=d\sigma/d\tau$, having a meaning independently of any coordinate system. Similarly in geometric language, the electromagnetic field is described by the second-rank, antisymmetric tensor $F$, which also requires no coordinates for its definition. This tensor produces a 4-force on any charged particle given by $\hat{P}_{\perp}\cdot(mdu/d\tau-eF\cdot u)=0$ [52]. This is the basic dynamics law for relativistic charged particles expressed in terms of geometric objects and automatically included the principle of relativity. The presence of the projector operator $\hat{P}_{\perp}$ suggests that we have only three independent equations. In the case of Maxwell’s equation we are able to rewrite the equations in the relativistic form without any change in the meaning at all, just with a change notations. It is important to noticed that the situation with dynamics equations is more complicated. In order fully to understand the meaning of the embedding of the dynamics law in the hyperplanes perpendicular to the world line, one must keep in mind that, above, we characterized the Newton’s equation in the Lorentz comoving frame as a phenomenological law. The system of coordinates in which the equations of Newton’s mechanics are valid can be defined as Lorentz rest frame. The relativistic generalization of the Newton’s second law to any Lorentz frame permits us to make correct predictions. The projector operator guarantees that this coordinate system restriction will be satisfied.
2. In general, the covariant equation of motion can be solved only by numerical methods; however, it is always attractive to find instances where exact solutions can be obtained. The simplest case of great practical importance is that the motion of a particle in a constant homogeneous magnetic field. From the solution of covariant equation of motion authors of textbooks [5, 52, 53] conclude that the covariant motion of a charge particle in a constant magnetic field is a uniform circular motion. The trajectory of the particle does not include relativistic kinematics effects and the Galilean vectorial law of addition of velocities is actually used. Actually, the old kinematics comes from the relation $d\tau=dt/\gamma$. This relation between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate in the lab frame, is based on the use of the absolute time convention. It is only after the authors of textbooks have make those replacement for $d\tau$ that authors obtain the usual formula for non-covariant trajectory.
3. The non-commutativity and non-associativity of the relativistic composition of non-collinear velocities was first recognized by Mocanu [54] and latter resolved by Ungar [55], by showing role played by the Wigner rotation in the transformation.
@@ 14 Relativity and Electrodynamics
The differential form of Maxwell’s equations describing electromagnetic phenomena in the Lorentz lab frame is given by Eq.(8). Now let us use these equations to discuss the phenomena called radiation. To evaluate radiation fields arising from an external sources in Eq. (8), we need to know the velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$. As discussed above, it is generally accepted that one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along non-covariant trajectory like $\vec{x}(t)$. The trajectory $\vec{x}(t)$, which follow from the solution of the corrected Newton’s second law Eq. (4) under the absolute time convention, does not include, however, relativistic kinematics effects.
We argue that this algorithm for solving usual Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the principle of relativity. However, the usual Maxwell’s equations in the lab frame, Eq. (8), are compatible only with covariant trajectories calculated by using Lorentz coordinates, therefore including such relativistic features as the relativity of simultaneity and the Wigner rotation.
We now consider the particle motion in a giving magnetic field. In their usual form, Maxwell’s equations are valid only in Lorentz reference frames. According to correct coupling of fields and particles, from the Eq. 10 follows that
@@
We find in this book that the Wigner rotation turns out to play an important role as the regulator of the velocity addition law. As we have already pointed, the four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame (1). The presentation of the time component simply as the relation $d\tau=dt^{\prime}/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
One of the consequences of non-commutativity of non-collinear Lorentz boosts is a difference between covariant and non covariant particle trajectories in a giving magnetic field. One can see that this essential point has never received attention by the physical community. As a result, a correction of the conventional radiation theory is required.
@@ 14.1 Why did the Error in Radiation Theory Remain so Long Undetected?
In order to answer this question, we shall discuss the subject more mathematically. For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. As example of such circumstance is a non-relativistic radiation setup. The non-relativistic asymptote provides the essential simplicity of the covariant calculation. The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories.
We want now to solve electrodynamics equations mathematically in a general way and consider the radiation associated with the succeeding terms in (multi-pole) expansion of the field in powers of the ratio $v/c$. Radiation theory is naturally developed in the space-frequency domain, as one is usually interested in radiation properties at a given position in space and at a certain frequency. In this book we define the relation between temporal and frequency domain via the following definition of Fourier transform pair:
@@
Suppose we are interested in the radiation generated by an electron and observed far away from it. In this case it is possible to find a relatively simple expression for the electric field (see Appendix I). We indicate the electron velocity in units of $c$ with $\vec{\beta}$, the electron trajectory in three dimensions with $\vec{r}(t)$ and the observation position with $\vec{r}_{0}$. Finally, we introduce the unit vector
@@
pointing from the retarded position of the electron to the observer. In the far zone, by definition, the unit vector $\vec{n}$ is nearly constant in time. If the position of the observer is far away enough from the charge, one can make the expansion
$\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)~{}.$ (30)
@@
where $\omega$ is the frequency, $(-e)$ is the negative electron charge and we make use of Gaussian units. A different constant of proportionality in Eq.(LABEL:revwied) and in the well known textbooks is to be ascribed to the use of different units and definition of the Fourier transform.
First we will limit our consideration to the case of sources moving in a non-relativistic fashion. According to the principle of relativity, usual Maxwell’s equations can always be used in any Lorentz frame where sources are at rest. The same considerations apply where sources are moving in non-relativistic manner. In particular, when oscillating, charge particles emit radiation, and in the non-relativistic case, when the velocities of oscillating charges $v\ll c$, dipole radiation will be generated and described with the help of the Maxwell’s equations in their usual form, Eq. (8).
Let’s examine in a more detail how the dipole radiation term comes about. The time $\vec{r}(t)\cdot(\vec{n}/c)$ in the integrands of the expression for the radiation field amplitude, Eq. (LABEL:revwied), can be neglected in the cases where the trajectory of the charge changes little during this time. It is easy to find the conditions for satisfying this requirement. Let us denote by $a$ the order of magnitude of the dimensions of the system. Then the time $\vec{r}(t)\cdot(\vec{n}/c)\sim a/c$. In order to ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$. This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges.
We consider the radiation associated with the first order term in the expansion of the Eq. (LABEL:revwied) in power of $\vec{r}(t)\cdot(\vec{n}/c)$. In doing so, we neglected all information about the electron trajectory $\vec{r}(t)$. In this dipole approximation the electron orbit scale is always much smaller than the radiation wavelength and Eq. (LABEL:revwied) gives fields very much like the instantaneous theory. So we are satisfied using the non-covariant approach when considering the dipole radiation theory.
But that is only the first and most practically important term. The other terms tell us that there are higher order corrections to the dipole radiation approximation. The calculation of this correction requires detailed information about the electron trajectory. Obviously, in order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach. However, correction to the multipole radiation is expected to be small. For example, covariant (trajectory) correction for quadrupole radiation is treated as a correction of higher (than quadrupole) order.

### v10

@@ 17.2 Derivation of the Covariant (BMT) Equation of Motion of Spin
To fully understand the meaning of embedding of the spin dynamics law in the Minkowski space-time, one must keep in mind that, above, we characterized the spin dynamics equation in the Lorentz comoving frame as a phenomenological law. The microscopic interpretation of the magnetic moment of a particle is not given. In other words, it is generally accepted that the spin dynamics law is a phenomenological law and the magnetic moment is introduced in an ad hoc manner. The system of coordinates in which the classical equations of motion for particle angular momentum are valid can be defined as Lorentz rest frame. The relativistic generalization of the three-dimensional equation $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ to any Lorentz frame permits us to make correct predictions.
17.3 Change Spin Variables
When Bargman, Michel, and Telegdi first discovered the correct laws of spin dynamics, they wrote a manifestly covariant equation in Minkowski space-time, Eq.(28), which describes the motion of the four-spin $S^{\mu}$. The derivation of this equation was very similar to the four-tensor equations that were already known to relativistic particle dynamics. The equation Eq.(28) is more complex than one might think. It is composed of a set of coupled differential equations. To find a solution directly from the system seems quite difficult, even for a very symmetric, uniform magnetic field setup. How to solve this four-tensor equation is an interesting question. In relativistic spin dynamics, it is done in one particular way, which is very convenient.
To apply Eq.(28) to specific problems it is convenient to introduce a three vector $\vec{s}~{}$ by the equation $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$. With the help of this relation, one can work out the equation of motion for $\vec{s}~{}$. Let us restrict our treatment of spinning particle dynamics to purely transverse magnetic fields. This means that the magnetic field vector $\vec{B}$ is oriented normally to the particle line motion. In this practically important case one has, after somewhat lengthy calculations [46]:
What must be recognized is that in the accepted covariant approach (indeed, Eq.(28) is manifestly covariant), the solution of the dynamics problem for the spin in the lab frame makes no reference to the three-dimensional velocity. In fact, the Eq.(29) includes relativistic factor $\gamma$ and vector $\vec{v}/c$, which are actually notations: $\gamma=\mathcal{E}/(mc^{2})$, $\vec{v}/c=\vec{p}c/\mathcal{E}$. All quantities $\mathcal{E},\vec{p},\vec{B}$ are measured in the lab frame and have exact objective meaning i.e. they are convention-independent. The evolution parameter $\tau$ is also measured in the lab frame and has exact objective meaning. For instance, it is not hard to demonstrate that $d\tau=mdl/|\vec{p}|$, where $dl$ is the differential of the path length.
Now we have an equation in the most convenient form to be solved. Suppose we let the charged spinning particle in the lab frame through a bending magnet with the length $dl$. We know that $d\theta=-eBdl/(|\vec{p}|c)$ is the orbital angle of the particle in the lab frame. Note that $d\tau=mdl/|\vec{p}|$. Then, Eq.(29) tells us that we may write the spin rotation angle with respect to the lab frame axes $\Omega d\tau$ as $\Omega d\tau=[(g/2-1)\gamma d\theta+d\theta]$.
The spin vector $\vec{s}$ directly gives the spin as perceived in a comoving system. If we say that in the lab frame, the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame, the spin makes this angle with the line motion of the lab frame. We must conclude that the conventional method used to explain the spin dynamics in the lab frame is very unusual. Indeed, the spin rotation measurement in the lab frame is interpreted from the viewpoint of the proper observer as viewing this of the lab observer.
17.4 An Alternative Approach to the BMT Theory
Above we described the BMT equation, Eq.(29), in the standard manner. It uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes. That means that we know the orientation of the proper spin with respect to the lab coordinate system which is moving with velocity $-\vec{v}$ and acceleration $-\gamma d\vec{v}/d\tau$ in the proper frame. Let’s look at what the equation Eq.(29) says in a little more detail. It will be more convenient if we rewrite this equation as
Now let’s see how we can write the right-hand side of Eq.(30). The first term is that we would expect the spin rotation due to a torque with respect to the proper frame axes $d\vec{\phi}_{L}=-eg\gamma\vec{B}d\tau/(2mc)=(g/2)\gamma d\vec{\theta}$. Here $d\vec{\theta}=-eBdl/(|\vec{p}|c)$ is the angle of the velocity rotation in the lab frame. It has also been made evident by our analysis in Section 8 that the angle of rotation $d\vec{\phi}_{W}=-e(\gamma-1)\vec{B}d\tau/(mc)=(\gamma-1)d\vec{\theta}$ corresponds to the Wigner rotation of the lab frame axes with respect to the proper frame axes. With these definitions, we have $d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=d\vec{\phi}_{L}\times\vec{s}-d\vec{\phi}_{W}\times\vec{s}$.
Now we introduce our new approach to the BMT theory, finding another way in which this complicated problem can be solved. We know that $d\vec{\phi}_{L}$ and $d\vec{\phi}_{W}$ are the rotations with respect to the proper frame axes. Actually, we only need to find the spin motion with respect to the lab frame axes. Now we must be careful about signs of rotations.
There is a good mnemonic rule to learn the signs of different rotations. The rule says, first, that the direction of the velocity rotation in the proper frame is the same as the direction of the velocity rotation in the lab frame. Second, the direction of the lab frame rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame. Third, the sign of the spin rotation due to a torque at $g>0$ (it is handy to remember that for an electron $g$ is positive and very nearly 2) means that the direction of the spin rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame.
Why the new derivation of the BMT equation is so simple? The reason is that the splitting of the particle spin motion with respect to the lab frame axes into the dynamic and kinematic parts can only be realized in the proper frame. In the proper frame, we do not need to know any more about a relativistic ”generalization” of the (phenomenological) classical equation of motion for the particle angular momentum. In this case, it is possible to separate the spin dynamics problem into the trivial (Larmor) dynamic problem and the kinematic problem of Wigner rotation of the lab frame axes in the proper frame.
17.5 Spin Tracking
Having written down the spin motion equation in a four-vector form, Eq.(28), and determined the components of the four-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation for the spin motion. This is a covariant relativistic generalization of the usual three-dimensional equation of magnetic moment motion, which is based on the particle’s proper time as the evolution parameter. We next wish to describe the spin motion with respect to the Lorentz lab frame using the lab time $t$ as the evolution parameter.
The frequency of spin precession can be written in the form $\varpi=\omega_{0}[1+\gamma(g/2-1)]$, where $\omega_{0}$ is the particle revolution frequency.
In the previous section, we saw that the particle path $\vec{x}(l)$ has an exact objective meaning i.e. it is convention-invariant. The spin orientation $\vec{s}$ at each point of the particle path $\vec{x}(l)$ has also exact objective meaning. In contrast to this, and consistently with the conventionality of the three-velocity, the function $\vec{s}(t)$ describing the spinning particle in the lab frame has no exact objective meaning.
which is based on the path length $l$ as the evolution parameter. These three equations correspond exactly to the equations for components of the proper spin vector that can be found from the non-covariant spin tracking equation Eq.(31). We want to emphasize that there are two different (covariant and noncovariant) approaches that produce the same spin orientation $\vec{s}(l)$ along the path. The point is that both approaches describe correctly the same physical reality and the orientation of the proper spin $\vec{s}$ at any point of particle path in the magnetic field has an objective meaning, i.e. is convention-invariant.
@@ 18 Relativity and Electrodynamics
The differential form of Maxwell’s equations describing electromagnetic phenomena in the Lorentz lab frame is given by Eq.(8). Now let us use these equations to discuss the phenomena called radiation. To evaluate radiation fields arising from external sources in Eq. (8), we need to know the velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$. As discussed above, it is generally accepted that one should solve the Maxwell’s equations in the lab frame with current and charge density created by particles moving along non-covariant trajectory like $\vec{x}(t)$. The trajectory $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law Eq. (4) under the absolute time convention, does not include, however, relativistic kinematics effects.
We now consider the particle motion in a given magnetic field. In their usual form, Maxwell’s equations are valid only in Lorentz reference frames. According to the correct coupling of fields and particles, from the Eq. 10 follows that
@@
[…]
@@ 18.1 Why did the Error in Radiation Theory Remain So Long Undetected?
To answer this question, we shall discuss the subject more mathematically. For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. An example of such a circumstance is a non-relativistic radiation setup. The non-relativistic asymptote provides the essential simplicity of the covariant calculation. The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories  [^63] Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation. The reason is that the ultrarelativistic assumption implies the paraxial approximation. Note that the geometry of the electron motion in a bending magnet has a cylindrical symmetry. There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation. It can be demonstrated that there is no need to use covariant particle tracking for derivation of the bending magnet radiation. However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field (see the next chapter for more details)..
@@
Suppose we are interested in the radiation generated by an electron and observed far away from it. In this case, it is possible to find a relatively simple expression for the electric field. We indicate the electron velocity in units of $c$ with $\vec{\beta}$, the electron trajectory in three dimensions with $\vec{r}(t)$ and the observation position with $\vec{r}_{0}$. Finally, we introduce the unit vector
@@
pointing from the retarded position of the electron to the observer. In the far zone, by definition, the unit vector $\vec{n}$ is nearly constant in time. If the position of the observer is far away enough from the charge, one can make the expansion $\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)$.
@@
where $\omega$ is the frequency, $(-e)$ is the negative electron charge and we make use of Gaussian units. A different constant of proportionality in Eq.(LABEL:revwied) and the well-known textbooks is to be ascribed to the use of different units and the definition of the Fourier transform.
Let’s examine in more detail how the dipole radiation term comes about. The time $\vec{r}(t)\cdot(\vec{n}/c)$ in the integrands of the expression for the radiation field amplitude, Eq. (LABEL:revwied), can be neglected in the cases where the trajectory of the charge changes little during this time. It is easy to find the conditions for satisfying this requirement. Let us denote by $a$ the order of magnitude of the dimensions of the system. Then the time $\vec{r}(t)\cdot(\vec{n}/c)\sim a/c$. To ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$. This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges.
We consider the radiation associated with the zero order term in the expansion of the Eq. (LABEL:revwied) in the power of $\vec{r}(t)\cdot(\vec{n}/c)$. In doing so, we neglected all information about the electron trajectory $\vec{r}(t)$. In this dipole approximation the electron orbit scale is always much smaller than the radiation wavelength and Eq. (LABEL:revwied) gives fields very much like the instantaneous theory. So we are satisfied using the non-covariant approach when considering the dipole radiation theory.
But that is only the first and most practically important term. The other terms tell us that there are higher-order corrections to the dipole radiation approximation. The calculation of this correction requires detailed information about the electron trajectory. To calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach. However, correction to the multipole radiation is expected to be small. For example, covariant (trajectory) correction for quadrupole radiation is treated as a correction of a higher (than quadrupole) order.